OpenStudy (anonymous):

State what the length of the other leg is in a right triangle whose hypotenuse is 13 and one leg is 12 by using the Pythagorean Theorem. A. 4 B. 5 C. 3 D. 6

OpenStudy (whpalmer4):

 By Pythagorean theorem, in a right triangle, the square of the length of the hypotenuse will be equal to the sum of the squares of the lengths of the two legs, or \[a^2+b^2 =c^2\] if \(c\) is the hypotenuse. You know that one of the legs is equal to 12, and the hypotenuse is equal to 13, so \[(12)^2+b^2=(13)^2\]or\[144 + b^2 = 169\]What is the value of \(b\) that makes that true?